OpenStudy (anonymous):

the equation of the circle with radius 3 and center (2,-5)

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2= r^2 center = (h,k) can you plug it in

OpenStudy (anonymous):

(x-2)^2 + (y+5)^2= 3^2